# Dense Retrievers Can Fail on Simple Queries: Revealing The Granularity Dilemma of Embeddings

### v1

@@ B.1 Query Types
  • Singleton Person: person-related queries, e.g. {CJK*}UTF8gkai男性 (male), {CJK*}UTF8gkai学生 (student).
  • Singleton Place: place-related queries, e.g. {CJK*}UTF8gkai健身房 (gym), {CJK*}UTF8gkai沙漠 (desert).
  • Singleton Object: other concrete entities, e.g. {CJK*}UTF8gkai食物 (food), {CJK*}UTF8gkai聊天记录 (chat history).
  • Singleton Concept: non-concrete concepts, e.g. {CJK*}UTF8gkai音乐 (music), {CJK*}UTF8gkai股票 (stocks).
  • Singleton Event: event/action-related queries, e.g. {CJK*}UTF8gkai婚礼 (wedding), {CJK*}UTF8gkai演唱会 (concert).
  • Conjunction: conjuncted entities, e.g. {CJK*}UTF8gkai烧烤加啤酒 (BBQ and beer), {CJK*}UTF8gkai樱花和传统建筑 (cherry blossoms and traditional building).
  • Simple Condition: entities/events with simple conditions, e.g. {CJK*}UTF8gkai演唱会相关群聊 (group chat regarding concerts), {CJK*}UTF8gkai睡觉的婴儿 (a sleeping baby).
  • Complex Condition: entities/events with more complex conditions, e.g. {CJK*}UTF8gkai一个人在田地里收割白菜 (a person harvesting cabbages in the field), {CJK*}UTF8gkai在沙发上的白猫 (a white cat next to a sofa).
@@ B.2 Label Distribution
As mentioned in Sec. 2, full labels are annotated for each query-passage pair in CapRetrieval, resulting in a total number of 1.3 million pair labels, comprising 4,683 positive pairs. The distribution of positive passages per query is provided in Fig. 2. Queries with the most number of relevant captions are general queries such as “{CJK*}UTF8gkai男性” (male), “{CJK*}UTF8gkai女性” (female), “{CJK*}UTF8gkai食物” (food).
@@ Appendix C Zero-Shot Evaluation
For embedding-based retrieval, all experiments follows the same setting: all text is firstly converted to lower case; all embeddings are normalized such that cosine similarity is the metric for retrieval. For the GTE-1.5B/7B, E5-7B and Qwen3 models, we use the instruction “Given an image search query, retrieve relevant image captions” in obtaining the query embedding for CapRetrieval.
For BM25, we use Jieba [^8] https://github.com/fxsjy/jieba for Chinese word segmentation, and use rank-bm25 for the BM25 python implementation.
@@ Appendix D Encoder Training
{CJK*}UTF8gkai炸鸡 (fried chicken) {CJK*}UTF8gkai一桌丰盛的餐点包括烤肉串、炸薯条和春卷。 (A table full of delicious dishes includes grilled meat skewers, French fries, and spring rolls.) 0 0.48
{CJK*}UTF8gkai图片展示了麦当劳麦辣鸡翅(2块)20次券的电子优惠券,售价185.3元, {CJK*}UTF8gkai单份低至7.9元。 (The image shows a digital coupon for 20 servings of McDonald’s Spicy Chicken Wings (2 pieces), priced at 185.3 RMB, bringing the cost as low as 7.9 RMB per serving.) 2 0.38
{CJK*}UTF8gkai紫色的花 (purple flower) {CJK*}UTF8gkai一辆紫色轿车停在路边,车顶和车窗上装饰有花束,车前挡风玻璃上有红色标签。 (A purple sedan is parked by the roadside, decorated with flower bouquets on the roof and windows, and a red tag on the front windshield.) 0 0.57
{CJK*}UTF8gkai图片中有四只紫色的蝴蝶,背景为浅紫色。 (The image features four purple butterflies against a light purple background.) 0 0.57
{CJK*}UTF8gkai一辆白色轿车停在树下,背景是紫色花田和远处的山脉。 (A white sedan is parked under a tree, with a purple flower field and distant mountains in the background.) 2 0.48
{CJK*}UTF8gkai图片展示了一片薰衣草田,背景是蓝天白云,文字内容为“只要你欢乐, {CJK*}UTF8gkai我就幸福浓浓。朋友,愿你欢乐无忧。早上好”。 (The image shows a lavender field with a background of blue sky and white clouds. The text reads: “As long as you’re happy, my heart is full of joy. My friend, may you be cheerful and carefree. Good morning.”) 2 0.37
@@
{CJK*}UTF8gkai西瓜 {CJK*}UTF8gkai图片中有一个装满水果的篮子,旁边有生菜、猕猴桃和小番茄。 0 0.50
{CJK*}UTF8gkai一辆装满西瓜的三轮车停在商店门口。 2 0.47
{CJK*}UTF8gkaiwatermelon In the picture, there is a basket full of fruit, with lettuce, kiwis, and cherry tomatoes next to it. 0 0.60
A tricycle loaded with watermelons is parked in front of the store. 2 0.55
Table 8: A more extreme example to illustrate the embedding granularity problem. For the query {CJK*}UTF8gkai西瓜 (watermelon) and passages in both Chinese and English accordingly, both the popular BGE large encoders bge-large-zh/en-v1.5 fail to retrieve the obviously more relevant passage (label 2) before the irrelevant one (label 0). Though this case can be simply resolved by lexical matching, it demonstrates that the embedding granularity problem exists across languages. In this work, we focus on the regarding evaluation in Chinese.

### v2

@@ B.1 Query Types
  • Singleton Person: person-related queries, e.g. 男性 (male), 学生 (student).
  • Singleton Place: place-related queries, e.g. 健身房 (gym), 沙漠 (desert).
  • Singleton Object: other concrete entities, e.g. 食物 (food), 聊天记录 (chat history).
  • Singleton Concept: non-concrete concepts, e.g. 音乐 (music), 股票 (stocks).
  • Singleton Event: event/action-related queries, e.g. 婚礼 (wedding), 演唱会 (concert).
  • Conjunction: conjuncted entities, e.g. 烧烤加啤酒 (BBQ and beer), 樱花和传统建筑 (cherry blossoms and traditional building).
  • Simple Condition: entities/events with simple conditions, e.g. 演唱会相关群聊 (group chat regarding concerts), 睡觉的婴儿 (a sleeping baby).
  • Complex Condition: entities/events with more complex conditions, e.g. 一个人在田地里收割白菜 (a person harvesting cabbages in the field), 在沙发上的白猫 (a white cat next to a sofa).
@@ B.2 Label Distribution
As mentioned in Sec. 2, full labels are annotated for each query-passage pair in CapRetrieval, resulting in a total number of 1.3 million pair labels, comprising 4,683 positive pairs. The distribution of positive passages per query is provided in Fig. 2. Queries with the most number of relevant captions are general queries such as “男性” (male), “女性” (female), “食物” (food).
@@ Appendix C Zero-Shot Evaluation
For embedding-based retrieval, all experiments follows the same setting: all text is firstly converted to lower case; all embeddings are normalized such that cosine similarity is the metric for retrieval. For the GTE-1.5B/7B, E5-7B and Qwen3 models that take a query instruction, we use the following instruction “Given an image search query, retrieve relevant image captions” in obtaining the query embedding for CapRetrieval.
Table 6 additionally provides zero-shot evaluation scores on the English version of CapRetrieval (dubbed CapRetrievalEn). GTE-7B achieves the best performance in both Chinese and English.
For BM25, we use Jieba [^8] https://github.com/fxsjy/jieba for Chinese word segmentation and NLTK [^9] https://www.nltk.org for English tokenization. rank-bm25101010https://pypi.org/project/rank-bm25 is used for the Python BM25 implementation.
nDCG@1 nDCG@5 nDCG@10
BM25 72.68 70.30 69.56
0.1B BGE 70.16 66.81 67.26
GTE 78.12 76.60 75.77
E5 77.72 75.52 74.53
0.3B BGE 63.53 61.89 61.94
E5 80.77 77.90 77.40
0.6B Qwen3 80.64 78.36 77.80
> 1B GTE-1.5B 73.21 72.46 72.04
GTE-7B 86.07 83.70 83.38
E5-7B 78.91 77.25 77.07
Qwen3-8B 80.11 78.83 78.38
Table 6: Evaluation results of zero-shot experiments on CapRetrievalEn (the according English version). The same multilingual encoder models are evaluated as in Table 2, except for BGE that is now switched to its dedicated English version.
@@ Appendix D Encoder Training
炸鸡
(fried
chicken)
一桌丰盛的餐点包括烤肉串、炸薯条和春卷。
(A table full of delicious dishes includes grilled meat skewers,
French fries, and spring rolls.)
0 0.48
图片展示了麦当劳麦辣鸡翅(2块)20次券的电子优惠券,售价185.3元,
单份低至7.9元。
(The image shows a digital coupon for 20 servings of McDonald’s Spicy Chicken Wings
(2 pieces), priced at 185.3 RMB, bringing the cost as low as 7.9 RMB per serving.)
2 0.38
紫色的花
(purple
flower)
一辆紫色轿车停在路边,车顶和车窗上装饰有花束,车前挡风玻璃上有红色标签。
(A purple sedan is parked by the roadside, decorated with flower bouquets on the roof
and windows, and a red tag on the front windshield.)
0 0.57
图片中有四只紫色的蝴蝶,背景为浅紫色。
(The image features four purple butterflies against a light purple background.)
0 0.57
一辆白色轿车停在树下,背景是紫色花田和远处的山脉。
(A white sedan is parked under a tree, with a purple flower field and
distant mountains in the background.)
2 0.48
图片展示了一片薰衣草田,背景是蓝天白云,文字内容为“只要你欢乐,
我就幸福浓浓。朋友,愿你欢乐无忧。早上好”。
(The image shows a lavender field with a background of blue sky and white clouds.
The text reads: “As long as you’re happy, my heart is full of joy. My friend,
may you be cheerful and carefree. Good morning.”)
2 0.37
@@
西瓜
图片中有一个装满水果的篮子,旁边有生菜、猕猴桃和小番茄。
0 0.50
一辆装满西瓜的三轮车停在商店门口。
2 0.47
watermelon
In the picture, there is a basket full of fruit, with lettuce, kiwis, and
cherry tomatoes next to it.
0 0.60
A tricycle loaded with watermelons is parked in front of the store.
2 0.55
Table 9: A more extreme example to illustrate the embedding granularity problem. For the query 西瓜 (watermelon) and passages in both Chinese and English accordingly, both the popular BGE large encoders bge-large-zh/en-v1.5 fail to retrieve the obviously more relevant passage (label 2) before the irrelevant one (label 0). Though this case can be simply resolved by lexical matching, it demonstrates that the embedding granularity problem exists across languages. In this work, we primarily focus on the regarding evaluation in Chinese.